OpenStudy (anonymous):

although things are in parenthesis if raised to the 0 power is the answer still 1?

OpenStudy (anonymous):

Yup :) one of the many rules of exponents!

OpenStudy (anonymous):

Anything to the power of 0 = 1

OpenStudy (anonymous):

okay perfect thanks guys ..

OpenStudy (anonymous):

(x^3y^2)^0 = 1 correct ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how do simplify questions like .. (x/5)^3

OpenStudy (australopithecus):

yes anything to the power of 0 is equal to 1 except 0^0

OpenStudy (australopithecus):

0^0 is undefined

OpenStudy (anonymous):

right that makes sense .

OpenStudy (australopithecus):

\[\frac{x^3}{5^3} = (\frac{x}{5})^3\]

OpenStudy (australopithecus):

that is a general rule btw

OpenStudy (australopithecus):

if both the numerator and denominator have the same exponent

OpenStudy (anonymous):

how do you know if they have the same exponent if the exponent is outside of the parenthesis

OpenStudy (australopithecus):

it is a general rule

OpenStudy (anonymous):

oh okay i see

OpenStudy (australopithecus):

\[\frac{x^a}{y^a} = (\frac{x}{y})^a\]

OpenStudy (anonymous):

okay . but how did you get the answer to the prior question , i'm a little confused

OpenStudy (anonymous):

why did you take the exponent away from the numerator and put it outside of the parenthesis and left the exponent on the denominator

OpenStudy (australopithecus):

I'm confused about what you are asking

OpenStudy (australopithecus):

sorry I mean x^3/125

OpenStudy (anonymous):

??

OpenStudy (anonymous):

(x/5)^3

OpenStudy (anonymous):

what do you do there

OpenStudy (anonymous):

 .. if that helps lol